The Cavo Company has an ROA of 10 percent, a profit margin of 12.75 percent, and an ROE of 18.75 percent.

What is the company’s total asset turnover? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Total asset turnover             times

What is the equity multiplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Equity multiplier             times
Answer #1
Given:
ROA 10% i.e 0.10
Profit margin 12.75% i.e 0.1275
ROE 18.75% i.e 0.1875
Total Assets Turnover :
To calculate Total Asset Turnover we will use ROA Formula:
ROA = Profit margin * Total Asset Turnover
0.10 = 0.1275*Total Asset Turnover
Total asset Turnover = 0.10/0.1275
0.784313725
Total Asset turnover = 0.78 Times
Equity Multiplier:
To find Equity Multiplier we will use ROE Formula:
ROE = ROA*Equity Multiplier
0.1875 = 0.10*Equity Multiplier
Equity Multiplier = 0.1875/0.10
= 0.1875/0.10
1.875
Equity Multiplier = 1.88 Times